OpenStudy (anonymous):

Ok here is another radical problem, let me type it out, ill be as fast as i can

OpenStudy (anonymous):

\[-3\sqrt[4]{x^{7}}-3\sqrt[4]{81x ^{7}}\]

OpenStudy (anonymous):

(-492xSeven to the seventh power)

jimthompson5910 (jim_thompson5910):

\[\Large -3\sqrt[4]{x^7}-3\sqrt[4]{81x^7}\] \[\Large -3\sqrt[4]{x^4*x^3}-3\sqrt[4]{81x^4*x^3}\] \[\Large -3\sqrt[4]{x^4}*\sqrt[4]{x^3}-3\sqrt[4]{81}*\sqrt[4]{x^4}*\sqrt[4]{x^3}\] \[\Large -3x\sqrt[4]{x^3}-3*3x*\sqrt[4]{x^3}\] \[\Large -3x\sqrt[4]{x^3}-9x*\sqrt[4]{x^3}\] \[\Large (-3x-9x)\sqrt[4]{x^3}\] \[\Large -12x\sqrt[4]{x^3}\] So \[\Large -3\sqrt[4]{x^7}-3\sqrt[4]{81x^7}=-12x\sqrt[4]{x^3}\]

OpenStudy (anonymous):

Thank you Jim :) I'm just so stupid :( I just cant figure these out.

jimthompson5910 (jim_thompson5910):

you'll get better with more practice

OpenStudy (anonymous):

I just have issues grasping it all, square roots just confuse me.

OpenStudy (anonymous):

Mine was right too.